Prev Next

SW&IT Meeting Minutes 2020-09-21

Attendees: Carl, David, Igor, Matthieu, Stephen, Xianyu


Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

LR20200907 - Bec, Matthieu 72.0
LR20200907 - Holmberg, Carl 80.0
LR20200907 - Edmeades, David 32.0
LR20200907 - Sola, Igor 40.0
LR20200907 - Hooper, Stephen 0.0
LR20200907 - Bryant, Trey 40.0
LR20200907 - Zhang, Yang 56.0
LR20200907 - Zhang, Xianyu 32.0

News from management

  • UA COVID-19 training status.
  • PEPSI / PETS and DR servers.

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports


  • Last week
    • OC, Management
    • ATT mostly
    • ROWG review Al's cleanup
  • This week
    • ROWG complete and distribute tech document
    • leadership workshop Tue/Wed 8-12am
    • ATT mostly


  • DIMM
    • LD moved it down from LLTH, gave him guidance on getting everything back together in old chassis. LD also checked out the DIMM enclosure and fiber as I was testing using the EDT tools before that, and that took him a good part of a day. I mention this, so we don't proceed blind.
    • We tried CentOS 6 in the old chassis, but didn't have any joy. LD put a new disk in for me, and also plugged in a USB stick.
    • I installed CentOS 4 and the old and LBTI EDT drivers. The "take" command no longer crashes the board (sometime last week I stopped crashing during "initcam", but I am not sure if I just imagined what was going on). I installed 64-bit CentOS 4, and the old EDT driver works, the LBTI and newest EDT drivers won't build for different reasons. I installed Fedora 30 and the newest EDT driver will not build (because of Fedora generic.h problem after F28 or 29). I installed CentOS 7. The newest EDT driver builds, and "take" works. I left the machine with both CentOS4 and CentOS 7 installs, defaulting to CentOS 7.
    • I pulled on the appropriate places to make the DIMM build on CentOS 7. I could have used Ice 3.4.2 work and Ice 3.4.2 from TCS, but I juat bumped it to 3.5, which is the oldeat that ia yum delivered. I will clean up compilation warnings and check the runtime over the next few daya.
    • We will try this out with the mount, and if it works, move it back into LLTH and try it out with the camera.
  • VPN work
  • robs and ROWG and 32-bit mountain stuff. Jenny, MODS, LUCI and OT.
  • UITS and wired port security.


  • ROWG technical
  • VPN work
  • DIMM continuation (handover, and beginning of week)
  • PEPSI recon, obs5, obs6.


  • Guido Timecapsule issues, ongoing  
  • Made text changes to safety type Issuetrak template, waiting for comments from DC
  • YZ VM upgrade on hold, waiting for him.
  • Synology remote access go-ahead, will set up this week
  • Pogge reverse SSH tunnel to mountain?
  • External VPN account policy ironed out
  • Cisco training
  • Archive permissions mostly sorted out, kicked to IA2 direct for fine tuning
  • Syno support
  • Office visit Monday evening
    • Drive swap
    • PEPSI server package for mountain transfer
  • I broke my Vault backup script, will fix it
  • Vault downtime announce for this weekend
  • ATT recording link destination?
  • Cisco train keeps going


  • Previous:
    • Some tcs-test rework for Alex Becker.
    • Some LBC and MODS misc discussions (LBC RB_Science program and 64-bit operations, MODS alert scripts etc.).
    • TMS work towards the next test (on Monday Sep 21).
    • Did the UA COVID-19 training.
  • Next:
    • TMS test on Monday (will probably miss the SWG meeting).
    • LBC 64-bit operation test discussion with Olga.
    • Need to work on AO-Cravat for most of the week.
    • Some issue tracks investigations if have time.


  • Last Week:
    • Some edits made to the structure and content of the investigation report
    • Some specific figures added to the report to make comparisons clearer (between different strategies used)
  • Next Week:
    • Finish up last things on report?
    • Next steps?


  • Last week:
    • AHS:
      • Fixed LBTI client byte to string bug, tested systemd script error.
      • Tested LBTI client during LBTI scheduled telescope time.
      • Debugged TMS laser client errors.
    • DMS: Reestablished iMonnit REST API access.
  • This week:
    • PR LBTI client, review test results with LBTI engineers.
    • Fill DMS iMonnit data gap.
    • Add filename change to RB_Science script, per JH.
    • Review DIMM environment recreation progress on newer Linux kernel.
    • Sit in on TMS and LBC discussions to catch up on project progress and direction.


  • Last week
    • Setup AdSec detailed information in the past 24 hours script for aoops
    • AdSec TSS with high wind analyze
    • Send/discuss AdSec list for ATT AdSec WG
    • ATT SW IT review
    • AO support
  • This week
    • Build a TCS simulator
    • Handle shell rested but dome open in AOS?
    • ATT SW list
    • Safety training
    • AO support


Last week:
  • LUCI
    • with the help of Chris Boylard did some testing on the Lab webio, more to be done.
  • PIT
    • restarted development of PIT Adming tool

This week:

  • Safety Training
  • PIT
    • Most of the week will be dedicated to implement the Close for Proposals opening and closing dates, and listing of current proposals per partner to the pit admin tool
  • LUCI
    • Possibly test with Geno lab webio.
    • Add more debugging information in Journalizer DB storing logic. Trying to figure out why sometimes the mask config file doesn't upload.
Topic revision: r15 - 21 Sep 2020, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback